solve this system of equations by substitution
2x+y=10
x=y+2

Answer by unlockmath(1688)

solve this system of equations by substitution
2x+y=10
x=y+2
Hello,
Here's what we do. Since one of the equations is already in the form of x = y + 2
we can plug y + 2 into the first equation. It would look like this:
2(y + 2)+y = 10 (Then distribute)
2y + 4 + y = 10 (combine terms)
3y + 4 = 10 (subtract 4 from both sides)
3y = 6 (Divide by 3)
y = 2 (now that we know that y = 2, we can find the value of x)
2x + 2 = 10 (subtract 2 from both sides)
2x = 8 (divide by 2 on each side)
x = 4 We got it solved. y = 2 and x = 4
Plug these into both equations and it should work out fine for you.
Good luck.
RJ Toftness
www.math-unlock.com